OpenStudy (anonymous):

I have already solved an equation ,The only thing i need help with is to graph it

OpenStudy (anonymous):

yes..?

OpenStudy (ledah):

you have the coordinates? or there are none

OpenStudy (anonymous):

Thanks Guys ..This was the question : Solve the system by graphing.If the system is inconsistent and has no solution,state this .If the system is dependent,write the form of the solution for any real number x. Below is my solution but having problem in graphing . 4x+2y=-8 y=(3/4)x+7 Now we rewrite as 4x+2y=-8 (3/4)x-y=-7 Divide the first equation by 2 and have 2x +y=-4 (3/4)x-y=-7 Add and get (2 3/4)x=-11 or (11/4)x=-11 or x=-4 and -16+2y=-8 gives y=4 (3/4)(-4)-4=-7

OpenStudy (ledah):

is the second one the x?

OpenStudy (ledah):

wait wait. this is confusing me..

OpenStudy (anonymous):

x=-11 and y=4